A watt is a unit of energy per unit time, and one watt (W) is equal to one joule per second (Jasa 1) A 60.0 W incandescent lightbulb produces about 5.00% of its energy as has an average wavelength of 5100 nm. calculate how many such photons are emitted per second by a 60.0W incandescent lightbulb.


1
Expert's answer
2022-03-06T15:16:13-0500

The amount of energy emitted as light per second is:


"E = 60J\times 5\% = 3J"

The energy in one photon is:


"E_1 = h\nu"

where "h = 6.63\times 10^{-34}J\cdot s" is the Planck's constant, and "\nu" is the frequency of the light. The frequecncy is given as follows:


"\nu = \dfrac{c}{\lambda}"

where "c = 3\times 10^8m/s" is the speed of light, and "\lambda = 5100nm = 5100\times 10^{-9}m" is the wavelength.

Thus, the number of the photons:


"N = \dfrac{E}{E_1} = \dfrac{E\lambda}{hc}\\
N = \dfrac{3J\cdot 5100\times 10^{-9}m}{6.63\times 10^{-34}J\cdot s\cdot 3\times 10^8m/s} \approx 7.69\times 10^{19}"

Answer. "7.69\times 10^{19}".